Young, Inc. 2018 Statement of comprehensive Income ($ in millions) Net sales $1384 Tess Cost of goods sold 605 Less: DepreciaA firm has a debt-equity ratio of 56. What is the total debt ratio? . Ο Ο Ο Ο
Answer #1

Debt to Equity ratio = 0.56

D/E = 0.56

E/D = 1/0.56

E/D = 1.78

E/D + 1 = 2.78

(E+D)/D = 2.78

Where E+D = Equity + Debt = Value = Assets

V/D = 2.78

D/V = 1/2.78 = 35.97 % = approx 36 % = Debt ratio